What is the density of a solid that has a mass of 136 g and a volume of 50 ml?

a. 0.36 g/ml b.  2.72 g/ml c. 2.77 g/ml d. 680 g/ml




1
Expert's answer
2021-04-07T10:08:29-0400
"\rho=\dfrac{m}{V}=\dfrac{136\ g}{50\ mL}=2.72\ \dfrac{g}{mL}."

Answer:

b. 2.72 g/mL